OpenStudy (anonymous):

Can someone help me solve this: differential equation// dy/dx= (e^3x) -3y

OpenStudy (anonymous):

\[\frac{ dy }{dy }+3y=e ^{3x}\] \[I.F=e ^{\int\limits 3 dx}=e ^{3x}\] solution is\[y*e ^{3x}=\int\limits e ^{3x}*e ^{3x}dx+c\] \[=\int\limits e ^{6x}dx+c\] =?